OpenStudy (anonymous):

Part a.) Solve the following equation and show your work. 2x + 2(x + 3) = 2(x + 7)

OpenStudy (anonymous):

@soty2013

OpenStudy (anonymous):

@some1

OpenStudy (anonymous):

ok so you have to distribute the 2 to (x+3) and the 2 on the right side to (x+7). It should look like this: 2x+2x+6=2x+14. Combine like terms in order to get this: 4x+6=2x+14. Subtract 2x from both sides. 2x+6=14. Subtract 6 from both sides. 2x=8. What is x equal to?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

shhh. i'm not supposed to give the answer, that's her job.

OpenStudy (anonymous):

lol thanks :)

OpenStudy (anonymous):

nppp :)